### v1

# RocqStar: Leveraging Similarity-driven Retrieval and Agentic Systems for Rocq generation

Abstract
Interactive Theorem Proving was repeatedly shown to be fruitful combined with Generative Artificial Intelligence. This paper assesses multiple approaches to Rocq generation and illuminates potential avenues for improvement. We highlight the importance of thorough premise selection for generating Rocq proofs and propose a novel approach, leveraging retrieval via a self-attentive embedder model. The evaluation of the designed approach shows up to 28% relative increase of the generator’s performance. We tackle the problem of writing Rocq proofs using a multi-stage agentic system, tailored for formal verification, and demonstrate its high effectiveness. We conduct an ablation study and show the use of multi-agent debate on the planning stage of proof synthesis.
1 Introduction
In recent years, the advent of Generative Artificial Intelligence (AI) has accelerated the process of developing new software. However, there are studies [21] showing that users who use AI assistants tend to introduce more bugs and vulnerabilities into their code, compared to those who write code on their own. Formal software verification could help mitigate the issue of bugs and security flaws, as it ensures that the software operates correctly and reliably in compliance with the given specification. Under the assumption of a well-formed specification, formal verification provides strong guarantees and an acceptance criterion for the generated code. Interactive Theorem Prover (ITP) is a software tool that assists the user with the development of formal specifications and proofs. To date, there exist several ITPs, such as Rocq (former Coq) [2], Lean [5], Agda [13], Isabelle [19], and others. Rocq is a mature ITP, which has experienced more than 30 years of continuous development and improvement. Rocq has an extensive track record of high-impact projects. For example, Rocq was used to verify the correctness of the CompCert C compiler [15], the only compiler, in which an extensive study found no bugs [33].
Verifying software has always been a rigorous and time-consuming process requiring much human effort. A number of solutions have been developed to help automate the process of theorem proving in Rocq. Proofs in Rocq are constructed from so-called tactics, which are elementary building blocks. Using tactics, the user manipulates the proof state — a data structure, which contains the current goal and the context of the proof. Thus, with every applied tactic, the task is transformed and could be solved recursively. Most solutions implement tactic-prediction approaches and employ beam search or a similar algorithm to navigate the search space. Tactician [4] is a KNN-based approach, which does similarity-based retrieval of tactics used in similar states. CoqGym [31] and Proverbot9001 [25] use Recurrent neural networks, Graph2Tac [24] proposed a novel graph-based neural tactic prediction. Thakur et al. [26] and Kozyrev et al. [14] instead build generation pipelines around general-purpose, cloud-hosted LLMs, so that no heavy computations occur on the user’s machine. CoqPilot [14], along with that, contributes a benchmarking framework and allows seamless integration of standalone tools into the workflow of Rocq’s user.
Many approaches call attention to the importance of premise selection, i.e., retrieving useful context information to advance generation. Yang et al. [32] introduced LeanDojo, a retrieval-augmented prover in Lean that significantly improves over non-retrieval baselines. Thompson et al. [27] present the Rango tool and report state-of-the-art performance on the CoqStoq benchmark, automatically synthesizing complete proofs for 32% of the theorems. The work highlights how strongly the well-formed context contributes to the success of Rango. Moreover, they show that proof retrieval is the most performant mechanism for premise selection. The proof retriever selects relevant previously completed proofs from the current project and provides them as references to the model. According to the evaluation, Rango proved 47% more theorems than the variant without a proof retriever. However, their mechanism of retrieving proofs relies on the baseline text similarity over states. In this work, we build on top of their research and propose a novel embedding model for Rocq statements. It is trained to predict the similarity of their proofs and shows relative improvement of up to 28% on the evaluation set.
Another promising direction in generative theorem proving that we have identified is Agentic Systems. Research by Kozyrev et al. [14] shows that current Rocq generation methods mostly struggle with complex reasoning tasks. Algorithms that perform proof search on top of a tactic generator slow down dramatically and suffer performance degradation as theorem complexity grows, due to the properties of tree-based search. Other neural methods, which apply LLMs, suffer from the same problem due to the inability of the model to handle complex reasoning tasks [11]. Agentic systems are known to address these problems; however, to our knowledge, there were close to no attempts to build an autonomous agentic system for an ITP. We build an extensive Model Context Protocol (MCP) server for Rocq and implement an autonomous Agentic System over it, utilizing various problem-specific solutions, such as multi-agent debate. We conduct an evaluation and show that our agentic system strongly outperforms all other previously benchmarked solutions in the CoqPilot’ work, raising the ratio of successfully proven theorems from 51% to 60%.
1.1 Contributions
The main contributions of this paper are:
RocqStar proof retriever  We propose a novel approach for premise selection in Rocq. Rocq suffers from the data-scarcity problem that is common to most ITPs. Aggregating the largest publicly available repositories, one could expect to collect roughly 300 million tokens of Rocq, and about the same for Lean. In contrast, open-source Python corpora easily exceed 100 billion tokens. To tackle this issue we contribute a convenient standalone tool BigRocq to extract additional data from Rocq code, utilizing the nature of Rocq’s system and the intermediate states of the proof. BigRocq bridges the gap between Automated Generation and Rocq’s ecosystem. Using BigRocq, we mine a dataset of 76,524 statements with corresponding proofs from 4 big projects and train a self-attentive embedder model, which learns to predict how close the proofs of given statements will be. In addition, we provide a pipeline to reproduce such embeddings for an arbitrary project, which offers even better results. We integrate the solution as a new retrieval approach for selecting context theorems in CoqPilot and evaluate it using CoqPilot’s benchmarking infrastructure. Compared to the baseline text similarity-based ranker, we show an improvement of 28% on the evaluation set. BigRocq tool, training dataset, and the code for training the embedder model are available at https://github.com/JetBrains-Research/rocqstar-rag. A model checkpoint is available at https://huggingface.co/JetBrains-Research/rocq-language-theorem-embeddings.
RocqStar agentic system  Addressing the lack of research of applying agentic systems to ITPs, we build an autonomous Agent for writing Rocq proofs. A custom MCP server built over coq-lsp [6] handles interaction with Rocq, its code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system. We implement an agentic system that includes such stages as planning, execution, and reflection. An ablation study demonstrates the critical role of planning, particularly the multi-agent debates (MAD) framework, in boosting performance. Evaluation shows that our end-to-end agent can solve 60% of theorems from the CoqPilot’s dataset. To deploy our AI Agent, we use privately available infrastructure called IDEFormer, but all our agent’s code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system/tree/main/rocqstar-agent.
The remainder of the paper is organized as follows. § 2 describes our Similarity-Driven Retrieval mechanism and § 3 introduces the agentic system. The retrieval component is evaluated in § 4.1 and the agent in § 4.2. § 4.3 provides an ablation study of the agentic system. We describe the related work in § 5 and conclude in § 6.
2 Similarity-driven Retrieval
A known problem in Retrieval Augmented Generation (RAG), applied to the domain of Interactive Theorem Proving (ITP), is premise selection [28, 9]. Premise selection is the task of retrieving facts from a given knowledge base to help the model advance the proof. Huang et al. [8] and Xu et al. [30] highlight the importance of a well-formed context, showcasing that the presence of irrelevant context information degrades the model’s performance.
We distinguish two ways of doing premise selection in Rocq. Hint selection — given a context $C$ and a tactic with an unknown positional argument, e.g. apply _, the task is to yield potential candidates for the argument. Proof selection, in turn, given theorem statement $S$, focuses on choosing other statements with their respective proofs, so that their presence in the context of the generation request would help the model with the generation of the proof for statement $S$. Most works [3, 12, 27, 32] on premise selection in Rocq and other ITPs focused on doing hint selection. However, Thompson et al. [27] and Kozyrev et al. [14] show that even a baseline proof selection significantly boosts the model’s capabilities and is stronger than hint selection. The baseline proof selection presented in both works [27, 14], given the target statement $s_{*}$ and a database of already proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, chooses theorems, statements of which have the maximum similarity to the target one. Similarity is defined by the BM-25 information retrieval technique [23] or Jaccard similarity index.
Both existing approaches suppose that if statements $s_{*}$ and $s_{i}$ are similar, their respective proofs $p_{*}$ and $p_{i}$ are similar as well:
$\operatorname{similarity(s_{*},s_{i})}\Longrightarrow\operatorname{similarity(p_{*},p_{i})}$
therefore assume that theorems $\{\left[s_{j},p_{j}\right]\}$, chosen in such a manner, are relevant while proving $s_{*}$. However, we show that this implication often does not hold. Let us define the proof similarity $D_{L}$ as the Levenshtein edit distance on lists of tactics, where the cost of substitution between two tactics is the Levenshtein distance over their strings. We include a Jaccard similarity term and add noise for robustness; otherwise, the proof-distance distribution over randomly selected pairs of theorems becomes U-shaped and the model fails to learn.
$p_{i}=\left[tac_{i_{0}},\dots,tac_{i_{m}}\right],\quad l_{i}=|s_{i}|,\quad D_{L}(p_{i},p_{j})=\frac{\mathrm{Lev}(p_{i},p_{j})}{\max(l_{i},l_{j})},\quad D_{J}(p_{i},p_{j})=1-\frac{|p_{i}\cap p_{j}|}{|p_{i}\cup p_{j}|}$
$\mathrm{proof\_distance}(p_{i},p_{j})=\alpha\,D_{L}(p_{i},p_{j})+(1-\alpha)\,D_{J}(p_{i},p_{j})+\gamma,\quad\alpha=0.7,\quad\gamma\sim\mathcal{U}(-\varepsilon,+\varepsilon)$
Considering 1,855,701 pairs of theorems from the IMM project [^1] IMM  https://github.com/weakmemory/imm, we compute correlations between statement similarities and respective proof similarities. In summary, BM25-based statement similarity shows a weak negative relationship with Levenshtein-based proof distance (Pearson $r$ = –0.154, Spearman $\rho$ = –0.171). In contrast, BM25-based proof similarity exhibits near-zero Pearson correlation ($r=0.029$) and a small positive Spearman correlation ($\rho=0.240$) — in both cases, effectively negligible.
To assess the issue of ineffective proof selection, we try to find such function $f(s_{i},s_{j})$ that correlates with $\operatorname{proof\_distance}(\operatorname{proof}_{i},\operatorname{proof}_{j})$ stronger than statement similarity. In this work, we introduce a neural method that learns vector embeddings for Rocq theorem statements, training them so that the distance between any two vectors mirrors the similarity between the theorems’ proofs.
2.1 Dataset mining
Along with other ITPs, Rocq struggles with data scarcity. To assess this issue, we mine additional data from the Rocq code. We utilize Rocq system’s functionality, preprocess theorems, and transform sequential proof structures into trees. Fig. 1 illustrates an example of the process. Since every node in such tree is a valid state, we can automatically construct a proof for it, recursively iterating through its subtree edges. By extracting the statements with corresponding proofs, we can enlarge an arbitrary dataset of Rocq theorems roughly by a magnitude of 4. Dataset format and its details are described in Appendix B.
Theorem test : forall n : nat, n = 0 \/ n <> 0. Proof. intros n. destruct n. - left; auto. - right; auto. Qed. ${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{2}}$ []${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{0}}$${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{1}}$${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{3}}$ []left; auto.intros n.destruct n.destruct n.right; auto. Lemma s1 (n : nat) : n = 0 \/ n <> 0. Proof. destruct n. - left; auto. - right; auto. Qed. Lemma s2 : 0 = 0 \/ 0 <> 0. Proof. left; auto. Qed. Lemma s3 (n : nat) : S n = 0 \/ S n <> 0. Proof. right; auto. Qed.
Figure 1: Processing theorems into trees. $s_{i}$ denotes a state
We call the proposed tool BigRocq and make it publicly available as a standalone component of our system. The idea of mining additional training data from the intermediate states of the ITP is not new; Kogkalidis et al. [12] conducted analogous research for the Agda [13] language. Similar research for Rocq also takes place; however, some of those works are highly dependent on the deprecated ways of communication with Rocq’s compiler [31] and do not support up-to-date versions of Rocq. In contrast, others implement similar ideas as a part of the training pipeline and do not allow for seamless reuse. Using BigRocq, we mine a total of 76,524 statements, collected from 344 files from 4 big Rocq projects.
2.2 Modeling
In our work, we formulate the problem as a self-supervised contrastive representation learning problem and train a self-attentive embedder model [18]. Given the dataset of pairs $\left[\operatorname{statement}_{i},\operatorname{proof}_{i}\right]$, and a similarity function, $\operatorname{f}(\operatorname{proof}_{i},\operatorname{proof}_{j})$, defined between two proofs, we try to learn such function $r$ (ranker), that takes corresponding Rocq statements as inputs, but behaves as close as possible to $f$. Given two statements, we learn to predict how similar their proofs shall be. In § 4, we evaluate the performance of the proposed model on the following task. Given a statement $s_{*}$ and a set of proven theorems, we want to choose $k$ premises and use them as context for generating a proof for $S$.
$\mathcal{T}=\left\{(p_{i},s_{i})\right\},\;\mathcal{S}$ =$=\left\{s_{i}\right\},\;r:\mathcal{T}\times\mathcal{S}\to\mathbb{R}$ $\operatorname{Top}_{k}(r,s_{*})$ $=\underset{(p_{i},s_{i})\in\mathcal{T}}{\arg\operatorname{top}_{k}}\,g(s_{i},s_{*})$
$\operatorname{Solve}(r,s_{*})$ $=\operatorname{Solve}\bigl{(}\operatorname{Top}_{k}(r,s_{*}),s_{*}\bigr{)}\in\{0,1\}$ $\mathcal{Q}(r)$ $=\mathbb{E}_{s_{*}\sim\mathcal{D}}\bigl{[}\operatorname{Solve}(r,s_{*})\bigr{]}$
Assume, without loss of generality, that by basic statement similarity we mean $BM25$-based similarity. As we have already shown in § 2, text similarity is a bad choice of $r$, which shows low correlation with the target function. However, it sets a strong baseline for our model. In practical applications, similar theorems occasionally have similar proofs. Accordingly, we have decided to fine-tune Microsoft’s 108-million-parameter encoder [7], originally pretrained on a combined corpus of programming and natural language texts. On a tiny extra test dataset, consisting of 50 theorems with corresponding hand-picked premises, raw CodeBert achieved an accuracy of 48%. This corresponds to roughly the same accuracy for ranking performed using the Jaccard-similarity metric on statements.
We train the model using InfoNCE [20] loss. In particular, given the statement $s$, on the dataset post-processing stage, we compute distances to other samples. We then mark a pair as positive if the distance between two proofs is less than a threshold $\tau_{pos}$, and we mark it as negative if the distance is greater than $\tau_{neg}$. Given the hyperparameter $k_{\operatorname{neg}}$ and sets of positive and negative pairs $P^{+}_{s}$ and $P^{-}_{s}$ we compute a per-statement loss term $\mathcal{L}_{s}$ as follows:
$\mathcal{L}_{s}=-\log\frac{\exp\bigl{(}\varphi(z_{s},z_{p})/T\bigr{)}}{\exp\bigl{(}\varphi(z_{s},z_{p})/T\bigr{)}+\displaystyle\sum_{j=1}^{k_{\operatorname{neg}}}\exp\bigl{(}\varphi(z_{s},z_{n_{j}})/T\bigr{)}}\qquad\bigl{(}p\in P^{+}_{s},\;n_{j}\in P^{-}_{s}\bigr{)}$
where $\varphi$ is a cosine similarity between $\ell_{2}$-normalized embeddings of statements. Experiments on the $k_{\operatorname{neg}}$ hyperparameter in our case showed little fluctuation in the results; however, $k_{\operatorname{neg}}=4$ procured the smoothest convergence, which aligns well with research by Wu et al. [29].
Given the particular shape of the sample distance distribution, during training we experienced the problem of the model converging too quickly on “easy” negatives — pairs, whose proofs (and typically their statements) are already far apart in the raw distance space. To keep informative gradients flowing, we add hard negative pairs; with some probability we treat a pair of statements as negative if $\tau_{\operatorname{hardneg}}\leqslant\operatorname{sim}(\text{proof}_{a},\text{proof}_{b})\leqslant\tau_{\operatorname{neg}}$. Introduction of negative samples helped to stabilize the training process; we have observed a less steep training curve and better generalization overall. Other training hyperparameters are listed in Appendix C.
3 Target theorem b : bool negb (negb b) = b. Premises Theorem plusn0: n : nat, (n + 1) =? 0 = false. Proof. ... Qed. Theorem add_comm: b c, andb b c = andb c b. Proof. ... Qed. $\vdots$ Theorem plus0: n : nat, 0 + n = n. Proof. ... Qed. 1
Pro LLM
Con LLM
Judge LLM
Initial planArgs againstArgs prom roundsPlanning Stage2Winning plan
Plan scoring
LLM
k plans
Critic LLM
Refer to caption Refer to caption
Executor LLM
analyzedAvailable toolsPlan 1: 0.95Plan k: 0.36Upd planif fails oftenEmbedderEmbedder
Figure 2: Agentic pipeline with RocqStar retriever.
3 Agentic System
Agent-based approaches are broadly used in code generation and repair tasks. Despite a large number of autonomous and semi-autonomous coding agents, they are not widely used in formal proofs generation and are not tailored to the Rocq specifics. To address this, we have implemented a RocqStar agentic system.
To allow interaction between the agent and Rocq’s system, we develop a REST API server that provides a set of tools that are useful during the execution. We apply our domain knowledge and construct these tools to bring an agent-driven proving process as close as possible to a human-driven one. Example of allowed function calls include checking validity of proofs, retrieving the valid prefix of given proof, gathering additional information about available entities in the context, and interacting with the context via performing commands like Print ?a. to identify the type of an argument or Search ?exp. to search for defined terms by a pattern. Toolset is described in detail in Appendix D. Interaction with Rocq’s system is carried out through its language server, coq-lsp [6]. To conform with a commonly used Model Context Protocol (MCP) and allow seamless agent interaction with the environment through tools, we implement an MCP server that wraps the REST API server. In the provided tool set, the most important is a proof-checking tool. It not only returns the answer whether the proof in question is valid, but in case of an erroneous proof, returns the error itself, where in the proof it happened, and the valid prefix before the error along with the remaining goals after this prefix. Such functionality enables the agent to keep track of the current proof state and benefit from partial proof progress.
3.1 Agent Logic
The input to the agent is presented as a target theorem without a proof and a file where it was declared, see box 31 of Fig. 2. Agent’s pipeline is logically split into two main stages: planning and execution. In the planning phase, multiple language models rigorously work out the strategy for the further implementation. During execution agents follow the plan aiming to generate the correct proof.
Planning Stage  We use the idea of multi-agent debates to produce a plan of how the agent should prove the given theorem. Specifically, we make two LLMs argue with each other about the plan: one of the LLMs produces the initial plan and defends it (pro LLM), the other one makes arguments against this plan (con LLM), see box 32 of Fig. 2. After several rounds of debates, the whole message history is sent to the judge LLM, that decides who is the winner of the debate, and returns the final plan. With this procedure, we generate $k$ plans. We send them to the plan scoring LLM and prompt it to assign a numerical score to each plan (the higher the better). After that, we select $l$ plans with the highest score and send them to the Execution Stage, see box 33 of Fig. 2.
Execution Stage  For each of the selected plans, we run an executor agent that follows the strategy, iteratively invoking the tools from the provided tool set — proof checker, context–inspection queries, search commands, and so on, as atomic actions. By calling these tools, it interacts with the environment via the MCP server. We track how many erroneous proofs were checked in a row, and if this number is higher than a fixed threshold (we set the threshold to five during evaluation), we call a critic model to evaluate the current progress of the proof and find the deviations from the selected plan. After that, we retrieve theorems along with their proofs, whose top-level goals are similar to the currently remaining goal, according to the cosine similarity between their RocqStar-ranker embeddings. We prompt the LLM to explain which tactic sequences could be helpful to finish our proof. We gather the generated criticism and send it to the replanner LLM to refine the current plan along with similar proofs and their analysis. The replanner is a separate language model that revises the plan based on the critic’s feedback and the retrieved examples. The whole message history is sent back to the executor agent. During the execution of each plan, $n$ tool calls are allowed. If valid proof is not found after $n$ tool calls, we denote the plan as failed. In this case, we ask a plan failure summarizer LLM to generate a short explanation of why the strategy execution failed and what happened during it. Then this summarized explanation is sent to the new execution stage with the next selected plan. This procedure is repeated until the correct proof is found or there are no more strategies to execute.
4 Evaluation
To evaluate our approach, partially and as a whole, we use the CoqPilot benchmarking framework. We required a dataset with a large number of human-written theorems and proofs. To compare our solution to existing ones, we decided to re-use the dataset by Kozyrev et al. [14]. It is limited to 300 theorems from the IMM Project [22], which was suitable for us in terms of computational and financial costs. The theorems are partitioned into three groups, corresponding to the difficulty level. The length (in tactics) of the human-written reference proof of the theorem estimates its difficulty. The sizes of each group are chosen with respect to the initial distribution of proof lengths in the project. Final group sizes and length ranges of each group could be found in Table 2. From now on, we will refer to the described dataset as the IMM-300 dataset. For smaller ablation studies we additionally prepared IMM-50, a 50-theorem subset of IMM, constructed with the same procedure. No theorems from the dataset were present in the training set of the RocqStar ranker embedding model. Moreover, the training set only contained partial theorem goals, no initial statements. Split of both datasets into groups, details, and limitations are described in Appendix A. Computational and financial resources used for experiments are described in Appendix F.
4.1 Retrieval Mechanism
We integrate our retrieval mechanism as a ranker into CoqPilot and evaluate it on the IMM-300 dataset with different models under the hood. We compare the performance of our ranker with the baseline approach, which works in the following manner. Given a target theorem statement $s_{*}$ and a set of proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, it ranks theorems in a descending order of $J(s_{*},s_{i})$, where $J(s_{*},s_{i})$ is Jaccard-similarity index and $S_{s_{i}}$ is a set of tokens inside a statement. The statement is split into tokens by whitespaces, commas, etc. Jaccard-similarity index is semantically almost the same as the BM-25 metric and produces the same numerical results. For each theorem in the dataset, we take theorems within the same file, sort them using the ranker (Jaccard or RocqStar, respectively), take the $k$ most relevant ones ($k$ is equal to 7 in our experiments), and send a request to the model to generate the completion. Chosen theorems are being sent as a few-shot prompt. Generation for each theorem is requested 12 times. If the Rocq’s system accepts any of the proofs, the theorem is considered solved. The target metric in our evaluation is the ratio of solved theorems. The evaluation results are presented in Table 1.
[…]
Ranker Jaccard RocqStar Jaccard RocqStar Jaccard RocqStar
GPT-4o 48% $\pm$ 5% 51%$\pm$ 5% 18% $\pm$ 4% 25%$\pm$ 3% 11% $\pm$ 4% 11% $\pm$ 5%
Claude 3.5 58% $\pm$ 5% 61%$\pm$ 4% 28% $\pm$ 5% 36%$\pm$ 5% 16% $\pm$ 5% 21%$\pm$ 5%
Table 1: Model performance under different ablations across all evaluation sets.
As can be seen from Table 1, our proposed RocqStar ranker outperforms the baseline Jaccard ranker on almost all experiments, showing reliable improvement. Most of the performance increase could be seen in the second group; we interpret these results as follows. For short theorems in the first group, the assumption that similar statements imply similar proofs often holds; therefore, both rankers perform close to one another. For complex theorems from the third group, it rarely happens that two theorems have significantly similar proofs, resulting in less advancement space for the model.
4.2 Agentic System
We evaluate our agentic system on the IMM-300 dataset, pursuing the goal to solve as many theorems as possible. For all of the parts of the planning stage, we use the Claude 3.5 Sonnet model, performing two rounds of debates between actors. Four plans are generated, and two are chosen for further execution. During execution, 20 tool calls are allowed from the MCP server. Additionally, after five proof-checking calls, the critic model (Claude 3.7 Sonnet) is invoked and analyzes whether a deviation from the initial plan has occurred. We use Claude 3.5 Sonnet for the execution and re-planning, and Google Gemini Flash 2.0 for other tasks, due to the necessity of a big context. Results of the evaluation are shown in Table 2.
Reference proof length $\leqslant 4$ 5 – 8 9 – 20 Total
Group size 131 98 71 300
OpenAI GPT-4o 50% 26% 15% 34%
OpenAI o1 66% 31% 8% 41%
Deepseek R1 58% 29% 11% 37%
Claude 3.5 Sonnet 73% 41% 27% 51%
LleMMa 7B 24% 11% 1% 15%
Tactician (synth) 45% 23% 10% 29%
RocqStar Agent 76% 56% 38% 60%
Table 2: Measuring the performance of different Rocq generation methods via CoqPilot
As shown in Table 2, our agentic system outperforms other benchmarked models inside the CoqPilot framework. The strongest model so far was Claude 3.5 Sonnet, which achieves 51% accuracy on the dataset, given 12 retries for each theorem. RocqStar agent achieves 60%, showing vigorous improvement. In terms of financial costs, we estimate a run of an agent on one theorem at 1.3 US dollars, compared to 0.25 US dollars for 12 requests to the pure Claude 3.5 Sonnet in CoqPilot.
4.3 Ablation study
Considering that software-verification tasks cannot be solved ad hoc, without explicit planning, we conduct an ablation study that measures how removing the Multi-Agent Debate (MAD) layer and reverting to single-pass planning affects the proportion of theorems successfully proved. In this experiment, we leave all other modules of the system unchanged, including the Plan Scoring LLM. We run two versions of an agent; the first generates plans via MAD, and the other generates plans by a single request to an LLM without further refinement. We evaluate both agents on the IMM-50 dataset and depict the results in Table 3. The same table depicts the performance of the Claude 3.5 Sonnet model on the IMM-50 dataset with 12 retries and RocqStar ranker; this result is provided for reference.
Reference proof length $\leqslant 4$ 5 – 8 9 – 20 Total
Group size 22 16 12 50
Agent with MAD 91% 56% 33% 66%
Agent w/o MAD 86% 44% 17% 56%
Claude 3.5 Sonnet 86% 37% 8% 52%
Table 3: Ablation study of Multi-Agent Debate at planning stage
Results in Table 3 confirm a consistent advantage for MAD across all three complexity groups, with the most significant improvement observed on harder theorems. This trend highlights the importance of MAD in solving composite multi-stage problems, such as complex proof construction. We attach an example of how MAD refines the execution plan and fixes it to manage to solve a previously unsolved theorem in Appendix E.
5 Related Work
Many Rocq generation methods improve generation using Retrieval Augmentation. Most of those works solve the hint selection problem, described in § 2. Those approaches build proofs tactic by tactic, retrieving relevant lemmas or definitions to use in the next step. The problem of searching for existing proofs that could advance the generation is barely described in the literature. CoqPilot [14] and Rango [27] pack the context for the generator model with theorems most similar to the one we are solving. Our work proposes a novel method of doing premise selection and shows improvement over the baseline from previous works [14, 27].
In our Multi-Agentic system, we distribute responsibility over different agents. Differentiating between models that handle natural reasoning and those that handle coding is common practice in agentic systems. The work of Li et al. [16] proposes a similar task force split into Thinker, Solver, Critic, and Debug agents. Liang et al. [17] introduces a Multi-Agent debate framework, shows that this approach encourages divergent thinking, and demonstrates its usability in complex reasoning tasks. We show that planning is essential for the formal verification pipeline. Theorem-proving demands a clear, high-level picture of the proof before executing any code. Running a multi-agent debate at the planning stage ensures rigorous evaluation of different approaches before interacting with Rocq’s system. We produce several plans for further execution. In a manner, close to Islam et al. [10], we assign scores to plans and run them in order of score decrease. To our knowledge, there were close to no attempts to building Agentic systems for ITPs. Yang et al. [32] have shown an initial proof of concept of an agent for Lean; however, their agent lacks automaticity, the pipeline incorporates only minimal tooling, and does not possess an explicit planning stage.
As a user interface, we utilize CoqPilot to integrate into the common Rocq’s programmer pipeline. CoqPilot is a VSCode [^2] VSCode: https://code.visualstudio.com plugin, facilitating access to Rocq generation methods for end-users.
6 Conclusion
We have presented a method to enhance retrieval-augmented generation in Rocq via leveraging neural premise selection using a self-attentive embedder model. We evaluated our proposed solution on a dataset of 300 Rocq theorems with two different generator models under the hood and showed a noticeable improvement of up to 28% relative to the baseline. Our result suggests that proof-aware premise selection considerably improves generation quality, particularly for medium-difficulty theorems, where the gap between statement similarity and proof similarity becomes more significant.
Our work pioneers the use of Agentic Systems applied to Formal Verification. We have implemented an advanced pipeline that includes rigorous planning via multi-agent debate, domain-specific tooling, and an adaptive executor–critic loop that iteratively refines proofs based on partial progress. We conclude that our RocqStar agent shows promising results, surpassing strong baselines and highlighting the applicability of agentic systems in the domain of theorem proving. We evaluated the idea of applying multi-agent debate to the planning stage of the agentic pipeline in our ablation study and demonstrated that it substantially improves the downstream proof success rate.
Acknowledgments and Disclosure of Funding
We thank Ekaterina Verbitskaia, Ivan Kabashnyi, Maksim Rozenberg, and Pavel Guliaev for their valuable comments regarding this work.
Appendix A IMM Evaluation Dataset
The collected IMM-300 dataset from the CoqPilot [14] work includes only thereoms with proofs of length no more than 20. For that reason the bucket with the most difficult theorems is labeled 9 — 20 tactics. This decision has been made, reflecting CoqPilot’s original focus on subgoals and shorter lemmas. Theorems of length no more than 20 tactics account for 83% of all proofs in the IMM project. As we take the same dataset, it possesses the same limitations. Therefore, we have not evaluated our solution on theorems, for which the reference proof contains more than 20 tactics. However, such theorems are quite rare.
The exact list of theorems used in each group could be found in the repository of the CoqPilot project: https://github.com/JetBrains-Research/coqpilot/blob/main/etc/docs/benchmark/.
A common problem with testing pipelines that include general-purpose LLM providers, such as OpenAI, is data contamination. We are aware, that the model could have possibly seen the human-written proofs, as the IMM project has been publicly available since a while. However, firstly, the model sees neither the theorem name, for which it is generating the proof, nor the proof goal exactly as it was in the initial file. As we treat them as proof states, rather than theorems, an LLM receives it in an equivalent, but slightly modified way. Secondly, as many of our experiments have shown, various quality of premise selection drasticly changes the behavior of the model. That hints that the model is not able to memorize all theorems and proofs. Lastly, data contamination issue was one of the things we had in mind, while developing BigRocq. One could pass a Rocq project into BigRocq as input, and for each theorem retrieve the sub-state, that is achieved after $k$ steps. On an example of $k=2$, the following theorem:
Lemma eq_trans (A : Type) : forall (x y z : A), x = y -> y = z -> x = z.
Proof.
intros x y z Hxy Hyz.
rewrite Hxy. (* State: (A : Type) (x y z: A) (Hxy: x = y) (Hyz: y = z) : y = z *)
rewrite Hyz.
reflexivity.
Qed.
Could be automatically tranformed into the following one:
Lemma eq_trans_modified (A : Type) (x y z: A) (Hxy: x = y) (Hyz: y = z) : y = z.
Proof.
rewrite Hyz.
reflexivity.
Qed.
The higher $k$ is chosen, the smaller would be the chances of data leakage, as the produced sub-state gets further and further from the original theorem.
Appendix B Encoder Training Dataset
One of the limitations of our BigRocq tool is that it cannot process theorems that contain so-called goal selectors. The following example illustrates how they work.
Theorem test2nat1 : forall n : nat, n = 0 \/ n <> 0.
Proof.
destruct n.
- left; auto.
- right; auto.
Qed.
This example could be rewritten with the use of goal selectors to the following proof:
Theorem test2nat1 : forall n : nat, n = 0 \/ n <> 0.
Proof.
intros n.
destruct n.
all: try (left; auto) || (right; auto).
Qed.
Due to the limited information we get from the Coq-LSP, our heuristic algorithm of transforming the proof into a tree breaks down. We cannot augment such theorems. The authors of CoqGym [31] also explicitly state that they do not handle theorems with goal selectors. They state that in their dataset, goal selectors occur in less than 1% data. The situation has changed since the work was published; the feature is now used more often but is still relatively rare. Goal selectors are an issue to be solved, and we are working on a solution by extracting some additional information from Rocq’s system through Coq-LSP.
The dataset is stored as a collection of JSON files and, due to its relatively small size, is stored within the repository, in the sub-directory with the model training code: https://github.com/JetBrains-Research/rocqstar-rag/tree/main/proof-embeddings/data/.
Dataset is split into training, validation, and test sets with proportions of 70%, 20%, and 10% respectively. Theorems from the same file do not appear in different sets. Parameters of building the dataset are listed in Table 4b. Pair of statements is considered as negative, if the proof_distance between them is greater than 0.65, and positive if it is less than 0.3. If the diatance is in range $[0.45,0.65]$, with probability of 30% it is also considered to be a negative pair (see hard negatives in § 2.2).
Appendix C Encoder Details
Parameter Value
algorithm AdamW (0.9, $0.9,\;0.99,\;\mathrm{e}{-2}$)
schedule linear warmup (10)
lr $4\mathrm{e}{-6}$
batch size (stmts) 32
dropout 0.1
(a) Optimization hyperparameters
Parameter Value
embedding dim 768
max sequence length 128
(positive, negative) threshold (0.3, 0.65)
threshold hard neg. 0.45
hard negatives prob. 30%
(b) Model&Dataset hyperparameters
Table 4: Hyperparameters of the embedder training
Hyperparameters used for training the embedder model are listed in Table 4. We have used microsoft/codebert-base as the base model and trained our embedder for 22000 steps with a batch size 32. We applied a dropout of 0.1 on the last layer of the model; the embedding dimension is 768, and the maximum sequence length is 128. We use AdamW optimizer with a linear warmup schedule for 10% of the training steps.
C.1 Visualizing RocqStar vs. Baseline Premise Selection
Here we try to illustrate the difference between different rankers and show an example of a theorem from IMM project, where our ranker outperforms the baseline. Figure 3 presents such an example.
[…]
Proof using.
[…]
 by rewrite H2.
Qed.
[…]
Proof using.
[…]
 lia.
Qed.
Figure 3: Theorems with dissimilar statements and similar proofs
If we measure the distance between theorems from Firgure 3 using the convential Jaccard distance, which is used by default in CoqPilot, we get $0.67$:
$\operatorname{Jaccard\_distance}(t1,t2)$ $=1-\frac{\left|\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{transitive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\cap\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{irreflexive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\right|}{\left|\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{transitive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\cup\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{irreflexive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\right|}$
$=1-\frac{1}{3}=0.67$
[…]
$\operatorname{proof\_sim}(t1,t2)$ $=0.32$
$\operatorname{embedder\_pred}(t1,t2)$ $=0.28$
Appendix D Agent toolset
Below are the tools that the agent uses to interact with Rocq’s system. The session is a utility abstraction, mainly handled by our middleware server under the MCP. It manages sessions and creates a new one when the agent starts proving a new theorem. Sessions are introduced to speed up type-checking and reduce overhead. When the session is started, we create a file, copy all required theorem’s context into it, type-check the context using Coq-LSP, and then start executing commands and continuously checking generated proofs in the context of this session.
  • list_coq_files: Returns a list of all Coq files in the project.
  • get_theorem_names: Retrieves available theorem names from a file, including the target theorem.
  • get_theorem_names_excl: Retrieves available theorem names from a file with target theorem excluded from the list.
  • get_current_target_state: Returns the stage of the proof for the target theorem in the current session.
  • get_theorem_with_proof: Given the theorem’s name, returns the theorem with its proof.
  • check_proof: Validates a proof (or a part of a proof) in the context of a session and returns either of the following:
    • (i)
      That there are no more goals to prove
    • (ii)
      Provided proof produces no errors, but the goal is not fully solved. Returns: updated goal state
    • (iii)
      The current goal is solved, but there are more goals at other depth levels. Returns: first unsolved goal at the closest depth level
    • (iv)
      Provided proof produces errors. Returns: error message
  • get_similar_proofs: Given theorem goal/statement as a string, it uses RocqStar ranker to retrieve theorems that are similar to the input statement and returns 15 most similar ones.
  • get_objects: Returns output of Rocq’s Print All command, issued in the context of the current session. This command prints all defined objects in the current file. In particular, that would mean printing all statements of theorems available above the one we are trying to prove at the moment of request.
  • about_term: Uses About Rocq’s Command in the current session. Accepts the term name as an argument. Outputs the term’s definition and a short description from the Coq-LSP.
  • search_pattern: Uses Rocq’s Search ?exp to search for a pattern in the current session’s file. An example of a valid command: Search (?a + ?b = ?b + ?a). It could be useful for finding lemmas that could be used in the proof.
  • print_term: Prints a term in the current session file. Uses Rocq’s Print command. Accepts the term name as an argument. Outputs the term’s definition.
  • check_term: Checks a term in the current session’s file. Uses Rocq’s Check command. Similar to print_term or about_term, but outputs only the type of the term. In the case of a theorem, it outputs its statement.
Appendix E Plan refinement via Multi-Agent Debate
We will illustrate the difference between the two plans; one was generated using a single call to a model, and the other was generated using multi-agent debate. The theorem we are trying to prove is the sub_IPROP from the src/imm/SubExecution.v file, that states:
Lemma sub_IPROP : subseteq (IPROP G’) (IPROP G).
The following plan was produced on the planning stage via a single request to the model, without multi-agent debate:
[…]
The proposed proof strategy is quite vague, too wordy, and imprecise. With this plan, the executor agent started by fetching similar theorems and then got stuck in a cycle of ad-hoc attempts to check incorrect proofs. The winning plan, on contra, was precise, concise, and clear. It had a top-level structure of the proof incorporated into the plan and made a winning solution to use the search command, which helped to finish the proof:
Proof.
unfold IPROP.
rewrite sub_W.
rewrite sub_is_ta_propagate_to_G.
done.
Qed.
Below, as a comparisson, is the plan, yielded by the multi-agent debate:
[…]
rewrite /IPROP /subseteq.
(* See expanded goal structure *)
intros x Hx.
(* Analyze Hx’s type carefully *)
(* Use discovered helper lemmas *)
3. Reasoning Steps: Work with explicit set inclusion based on subseteq definition Use any discovered lemmas about G/G’ relationship Preserve flexibility to adapt based on actual definitions
[…]
Appendix F Experiments compute resources
Our experiments in the evaluation part mainly use cloud LLM providers and therefore require minimum compute, but comprehensive financial resources.
Embedder training  During our final training run of the embedding model for RocqStar ranker, the model consumed roughly 43 GB of GPU-process memory and only about 6% of the host’s RAM. Over 13.5 hours on a single NVIDIA H100 accelerator (with 20 CPU cores and 160 GB of system memory), disk usage grew steadily from 28 GB to 76 GB as checkpoints and logs accumulated. GPU utilization stabilized above 85% shortly after the warmup phase and remained near saturation for the remainder of training. To sum up, our setup runs comfortably on a single GPU node with modest additional CPU and memory overhead.
Embedder model evaluation  Experiments were conducted on a single MacBook Pro with an M1 chip. The only computationally expensive part of the experiments (for the local machine) is launching multiple Coq-LSP servers at once (CoqPilot benchmark does that to optimize the time of the experiments and accelerate type-checking). As we use a middleware service over LLM APIs, our financial estimations might not be accurate. However, we roughly estimate 12 generation attempts per theorem with seven contextual theorems at 12 cents per theorem for Claude 3.5 and 7 cents for GPT-4o. We run an experiment for 300 theorems and repeat it three times, resulting in 114 US Dollars.
Agent evaluation  In the case of the agent evaluation, we ran the experiment only once and did not provide the confidence intervals due to financial limitations. We ran our agent on the IMM-300 dataset, and afterward, we compared two versions of the agent on the IMM-50 dataset in our § 4.3. That results in 400 attempts to prove different theorems. We estimate a single attempt at 1.3 US dollars. Therefore, we estimate the cost of the evaluation of the agent to be 520 US dollars.